### v1

# ORBGRAND Is Exactly Capacity-achieving via Rank Companding

Abstract
Among guessing random additive noise decoding (GRAND) algorithms, ordered reliability bits GRAND (ORBGRAND) has attracted considerable attention due to its efficient use of soft information and suitability for hardware implementation. It has also been shown that ORBGRAND achieves a rate very close to the capacity over additive white Gaussian noise channels with antipodal inputs. In this work, it is further established that, via suitably companding the ranks in ORBGRAND according to the inverse cumulative distribution function (CDF) of channel reliability, the resulting CDF-ORBGRAND algorithm exactly achieves the mutual information of general binary-input memoryless channels under symmetric input distribution, i.e., the symmetric capacity. This result is then applied to bit-interleaved coded modulation (BICM) systems to handle high-order input constellations. Via considering the effects of mismatched decoding due to both BICM and ORBGRAND, it is shown that CDF-ORBGRAND is capable of achieving the BICM capacity, which was initially derived by treating BICM as a set of independent parallel channels.
I Introduction
Guessing random additive noise decoding (GRAND) [1, 2] has recently attracted attention as a new approach to decoding in next-generation communication scenarios. GRAND operates by progressively guessing the error pattern (EP) that the transmitted codeword has experienced through the channel. This approach centering around EPs makes the GRAND paradigm well suited for codes in the high-rate high-reliability regime.
GRAND was originally designed for hard-decision channels. This is also reflected in its name: only when the noise is discrete it is meaningful to “guess” it. For more general channels, such as additive white Gaussian noise (AWGN) channels, incorporating channel soft information can improve decoding performance [3, Ch. 10], and we need to guess the EP rather than the noise itself. Symbol reliability GRAND (SRGRAND) [4] uses a single-bit reliability indicator to mark whether a channel output symbol is reliable. Soft GRAND (SGRAND)[5] directly exploits the exact log-likelihood ratio (LLR) magnitudes of the channel output to determine the sequence of EPs for guessing, thereby fully leveraging the soft information and achieving maximum-likelihood (ML) decoding. However, the requirement of exact LLR magnitudes in SGRAND necessitates on-the-fly generation of EPs, posing challenges for hardware implementation. Another variant of GRAND, ordered reliability bits GRAND (ORBGRAND) [6] generates EPs based on the ranking of LLR magnitudes rather than their exact values. This approach enables offline generation of EPs and streamlines hardware implementation [7, 8, 9]. It has also been shown in [10] through an information-theoretic analysis that for AWGN channels with antipodal inputs, the rate achieved by ORBGRAND is very close to the channel mutual information, i.e., the capacity in that setup. There also exist several works that treat GRAND for more general channels such as fading channels [11, 12, 13, 14, 15, 16].
In this paper, we consider a modification of ORBGRAND, which essentially adopts a companding technique to transform the rank of each LLR magnitude according to its inverse cumulative distribution function (CDF). The resulting algorithm is termed CDF-ORBGRAND, whose basic idea has initially appeared in [6] and formal description provided in [17, 18]. Via such a companding technique, the decoding metric is rendered more aligned with the channel law, while still maintaining its amenability to hardware implementation. We analyze the achievable rate of CDF-ORBGRAND for general binary-input memoryless channels under symmetric input distribution, and prove via a random coding argument akin to that in [10] that, CDF-ORBGRAND exactly achieves the mutual information, i.e., the symmetric capacity, without any rate loss. In contrast, for fading channels, ORBGRAND without rank companding may incur a significant rate loss compared to the symmetric capacity; see Fig. 1 in Section III-A.
On the basis of CDF-ORBGRAND for binary-input channels, we proceed to the investigation of bit-interleaved coded modulation (BICM) systems [19, 20]. This investigation is motivated by the prevalence of high-order modulation in modern communication systems, wherein BICM has been an extensively used technique in fading channels. By jointly considering BICM and GRAND as a mismatched decoding rule, we derive achievable rates of both CDF-ORBGRAND and ORBGRAND for BICM. Remarkably, we show that CDF-ORBGRAND achieves the BICM capacity, which was initially derived in the literature by treating BICM as a set of independent parallel channels [19], and subsequently established by typicality decoding or by mismatched decoding with appropriately chosen metrics [21]. The BICM capacity-achieving property of CDF-ORBGRAND indicates that the ranking companding technique is effective even for handling the heterogeneous bit channels corresponding to different positions of the labeling of constellation symbols. In contrast, without rank companding, ORBGRAND incurs a substantial rate loss which becomes even more pronounced as modulation order increases, as confirmed by numerical results.
The remaining part of this paper is organized as follows: Section II introduces the system model, along with GRAND and its variants for general binary-input memoryless channels. Section III proves that CDF-ORBGRAND exactly achieves the symmetric capacity. Section IV considers a BICM system model, and applies GRAND therein. It further proves that CDF-ORBGRAND achieves the BICM capacity. Section V concludes this paper. Technical proofs are provided in Appendices.
II Preliminaries
II-A System Model
Our basic channel model is a general binary-input memoryless channel, which, without loss of generality, is of input alphabet $\{+1,-1\}$, with the output probability density function (pdf) denoted by $q^{+}(y)$ when input $x=+1$, and by $q^{-}(y)$ when input x=1x=-1. Here $q^{+}(y)$ and $q^{-}(y)$ are general, without assuming any symmetry.
A codebook has length $N$ and rate $R$ nats per channel use, corresponding to $\lceil e^{NR}\rceil$ possible messages. When transmitting message $w$, the corresponding codeword is denoted by $\underline{x}(w)=[x_{1}(w),\cdots,x_{N}(w)]$. In this paper we conduct a random coding analysis, and hence assume that the elements of $\underline{x}(w)$ are independent and identically distributed (i.i.d.), uniformly drawn from $\{+1,-1\}$. A message $\mathsf{W}$ is uniformly selected from $\{1,\ldots,\lceil e^{NR}\rceil\}$ for transmission. Given the channel output vector $\underline{\mathsf{Y}}=[\mathsf{Y}_{1},\cdots,\mathsf{Y}_{N}]$, we define the LLRs as
$\mathsf{T}_{i}=\ln\frac{q^{+}(\mathsf{Y}_{i})}{q^{-}(\mathsf{Y}_{i})},\quad i=1,\cdots,N,$
and call their magnitudes, $\lvert\mathsf{T}_{1}\rvert,\cdots,\lvert\mathsf{T}_{N}\rvert$], the channel reliability vector. The CDF of $\lvert\mathsf{T}\rvert$ is denoted by $\Psi(t)$, for $t\geq 0$. We also introduce the sets $\mathcal{R}_{1}=\left\{y\lvert q^{+}(y)<q^{-}(y)\right\}$ and $\mathcal{R}_{2}=\left\{y\lvert q^{+}(y)>q^{-}(y)\right\}$.
For $i=1,\cdots,N$, we denote $\mathsf{R}_{i}$ as the rank of $\lvert\mathsf{T}_{i}\rvert$ among the channel reliability vector, from $1$ (the smallest) to $N$ (the largest).
For technical reasons, we impose a few additional assumptions on the behavior of $q^{+}(y)$ and $q^{-}(y)$, as follows.
Assumption 1
Both $q^{+}(y)$ and $q^{-}(y)$ are light-tailed, in the sense that there exists a sufficiently large constant $M_{1}>0$, a constant $a>0$ and a polynomial $S_{1}(\cdot)$ satisfying
$q^{\pm}(y)<S_{1}(\lvert y\rvert)e^{-a\lvert y\rvert},\quad\forall\lvert y\rvert>M_{1}.$ (1)
Assumption 2
The growth of $\lvert t\rvert=\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert$ is polynomially bounded; that is, there exists a sufficiently large constant $M_{2}>0$ and a polynomial $S_{2}(\cdot)$ satisfying $\lvert t\rvert<S_{2}(\lvert y\rvert),\forall\lvert y\rvert>M_{2}$.
Assumption 3
The CDF of $\lvert\mathsf{T}\rvert$, $\Psi(\cdot)$, and its inverse, $\Psi^{-1}(\cdot)$, are smooth with finite first, second, and third derivatives.
II-B GRAND and Its Variants
Upon receiving a channel output vector $\underline{y}$, we compute the channel reliability vector $\left[\lvert t_{1}\rvert,\cdots,\lvert t_{N}\rvert\right$] and the corresponding hard-decision vector $\underline{x}_{\text{hard}}=[\text{sgn}(t_{1}),\cdots,\text{sgn}(t_{N})]$, where $t_{i}=\ln\frac{q^{+}(y_{i})}{q^{-}(y_{i})}$ is the LLR and $\text{sgn}(t)=1$ if $t\geq 0$ and 1-1 otherwise.
In general, we can collectively denote the EPs used by GRAND as a $2^{N}\times N$ $\pm 1$-valued matrix $P$: in the $q$-th query, the $q$-th row of $P$ is used as the EP for testing, such that if $P_{q,i}=-1$ the sign of the $i$-th entry of $\underline{x}_{\text{hard}}$ is flipped, and otherwise it remains unchanged. If $\underline{x}_{\text{hard}}$ after flipping is a codeword, it is declared as the decoder output; otherwise, if all rows of $P$ are exhausted without finding a codeword, a decoding failure is declared.
For ORBGRAND, we arrange the rows of $P$ so that the sum reliability of the $q$-th row, $\sum_{i:P_{q,i}=-1}r_{i}$, is non-decreasing with $q$, where $r_{i}$ is the rank of $\lvert t_{i}\rvert$ introduced in the previous subsection. Efficient algorithms for generating such $P$ are available (see, e.g., [6, 7]). Since $2^{N}$ is typically an exceedingly large quantity, in practice $P$ is usually truncated to its first $Q$ rows, with $Q$ being the maximum number of queries allowed.
As shown in [10, Sec. II], ORBGRAND and many other variants of GRAND can be represented by the following unified decoding rule, as long as no truncation of $P$ is applied, i.e., $Q=2^{N}$:
$\hat{w}=\mathop{\arg\min}\limits_{w=1,\cdots,\lceil e^{NR}\rceil}\frac{1}{N}\sum_{i=1}^{N}\gamma_{i}(\underline{y})\cdot\mathbf{1}(\text{sgn}(t_{i})\cdot x_{i}(w)<0).$ (2)
Different choices of $\left\{\gamma_{i}\right\}_{i=1,\cdots,N}$ correspond to different choices of $P$ and thus different variants of GRAND:
  • When $\gamma_{i}(\underline{y})=1$, (2) is the original GRAND [1];
  • When $\gamma_{i}(\underline{y})=\left\lvert t_{i}\right\rvert$, (2) is SGRAND [5], which is equivalent to ML decoding;
  • When $\gamma_{i}(\underline{y})=\frac{r_{i}}{N}$, (2) is ORBGRAND [6];
  • When $\gamma_{i}(\underline{y})=\Psi^{-1}\left(\frac{r_{i}}{N+1}\right)$, where $\Psi^{-1}$ denotes the inverse function of the CDF of $\lvert\mathsf{T}\rvert$, $\Psi$, [^1] In statistics, $\Psi^{-1}$ is also called the quantile function of $\lvert\mathsf{T}\rvert$. (2) is called CDF-ORBGRAND.
The form of (2) facilitates information-theoretic analysis, as will be conducted in the following sections. A decoding error event occurs if $\hat{\mathsf{W}}\neq\mathsf{W}$, and we define the achievable rate of GRAND as follows.
Definition 1
For the system model in Section II-A, a rate $R$ is achievable under GRAND with given $\{\gamma_{i}\}_{i=1,\ldots,N}$ if there exists a sequence of codebooks of length $N=1,2,\ldots$, such that the average error probability of decoding for the decoding rule (2) asymptotically vanishes as $N$ grows without bound.
If we do not restrict the decoding rule to be (2) and apply ML decoding, then according to Shannon’s channel coding theorem, the supremum of all achievable rates is the mutual information $I(\mathsf{X};\mathsf{Y})$ with $\mathsf{X}$ uniformly distributed over $\{+1,-1\}$. This mutual information thus serves as the ultimate performance limit for the considered system model, and is termed the symmetric capacity in the sequel.
III Capacity-achieving Property of CDF-ORBGRAND
To motivate CDF-ORBGRAND, it is helpful to inspect the decoding rule (2). Based on the convergence of empirical distributions, as $N$ gets large, the distribution of $\mathsf{R}_{i}/(N+1)$ gets close to that of $|\mathsf{T}_{i}|$, i.e., $\Psi$. [^2] This observation has been used in [10] to explain the behavior of ORBGRAND, by introducing a variant of GRAND termed cdf-GRAND therein. Unfortunately, cdf-GRAND relies on the exact channel reliability vector, and hence is not amenable to hardware implementation. Thus, by conducting the inverse operation, i.e., $\Psi^{-1}$, over $\mathsf{R}_{i}/(N+1)$, we essentially return to $|\mathsf{T}_{i}|$, and this is exactly what CDF-ORBGRAND does. Due to this intuitive argument, CDF-ORBGRAND should behave similarly to SGRAND, and this assertion will be formalized and established in this section. Note that in CDF-ORBGRAND since $\left\{\gamma_{i}\right\}_{i=1,\cdots,N}$ simply compands the rankings $\left\{\mathsf{R}_{i}\right\}_{i=1,\cdots,N}$, which can be computed and stored offline, its implementation is still similar to that of ORBGRAND, amenable to parallelism [22].
III-A Main Result
Neither ORBGRAND nor CDF-ORBGRAND is ML decoding, and furthermore, their metrics, $\left\{\gamma_{i}\right\}_{i=1,\cdots,N}$, are $N$ correlated random variables due to ranking, we cannot directly invoke existing information-theoretic formulas for evaluating their achievable rates. In this paper, we essentially establish the achievable rate of CDF-ORBGRAND as the generalized mutual information (GMI) of the considered decoding rule, following the approach of mismatched decoding (see, e.g., [23, 24]). As revealed by our result below, although CDF-ORBGRAND is inherently mismatched, its achievable rate surprisingly coincides with the symmetric capacity. In other words, from an information-theoretic perspective, no rate loss is incurred when applying CDF-ORBGRAND.
Theorem 1
For the system model in Section II-A, CDF-ORBGRAND achieves the symmetric capacity, i.e.,
$I(\mathsf{X};\mathsf{Y})$ $=\ln 2-\frac{1}{2}\int_{-\infty}^{\infty}\ln\left(1+\frac{q^{-}(y)}{q^{+}(y)}\right)q^{+}(y)\mathrm{d}y$
$\hskip 30.00005pt-\frac{1}{2}\int_{-\infty}^{\infty}\ln\left(1+\frac{q^{+}(y)}{q^{-}(y)}\right)q^{-}(y)\mathrm{d}y$ (3)
in nats/channel use.
Proof:
See Section III-B. ∎
In our previous work[25], we derived the GMI of ORBGRAND for the system model in Section II-A as
$I_{\text{ORB}}$ $=\ln 2-\inf\limits_{\theta<0}\Bigg\{\int_{0}^{1}\ln(1+e^{\theta t})\mathrm{d}t$
$\hskip 40.00006pt-\theta\cdot\frac{1}{2}\int_{\mathcal{R}_{1}}\Psi\Bigl(\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert\Bigr)q^{+}(y)\mathrm{d}y$
$\hskip 40.00006pt-\theta\cdot\frac{1}{2}\int_{\mathcal{R}_{2}}\Psi\Bigl(\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert\Bigr)q^{-}(y)\mathrm{d}y\Bigg\}$ (4)
in nats/channel use. According to the proof of Theorem 1 in Section III-B, the symmetric capacity (3) can be equivalently rewritten as
[…]
$\quad-\theta\cdot\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad-\theta\cdot\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y\Bigg\}$ (5)
[…]
We conduct some numerical experiments to compare (4) and (5), for antipodal modulation over AWGN, additive white Laplacian noise (AWLN), and Rayleigh fading channels with receiver-side perfect channel state information (CSI). These channels can be written as
$\mathsf{Y}=\sqrt{P}\mathsf{H}\mathsf{X}+\mathsf{Z},$
where $\mathsf{H}$ is the channel coefficient and $\mathsf{Z}$ is the additive noise. The corresponding LLR $\mathsf{T}$ is:
  • AWGN channel, $\mathsf{H}=1$ and $\mathsf{Z}\sim\mathcal{N}(0,1)$: $\mathsf{T}=2\sqrt{P}\mathsf{Y}$.
[…]
  • Rayleigh fading channel with receiver-side perfect CSI, $\mathsf{H}\sim\mathcal{CN}(0,1)$ and $\mathsf{Z}\sim\mathcal{CN}(0,1)$: $\mathsf{T}=4\sqrt{P}\Re(\mathsf{H}^{*}\mathsf{Y})$. We point out that here $q^{\pm}(\cdot)$ should be evaluated for the augmented channel output $(y,h)$, i.e., including the receiver-side perfect CSI. [^4] We clarify that in the numerical experiments in [25], the GMI of ORBGRAND for Rayleigh fading channels with receiver-side perfect CSI was incorrectly computed, due to an error in the program evaluating the integrals in (4). In the current paper, we rectify the error and provide updated numerical results.
The curves of $I_{\text{ORB}}$ and $I_{\text{CDF-ORB}}$ for different signal-to-noise ratio (SNR) values are displayed in Fig. 1. Their comparison indicates that in AWGN channel, ORBGRAND almost achieves the symmetric capacity, a result already established in [10]; that in AWLN channel, a noticeable gap appears; and that for Rayleigh fading channel, the gap becomes even more pronounced.
Figure 1: Comparison between $I_{\text{ORB}}$ and $I_{\text{CDF-ORB}}$ under AWGN, AWLN, and Rayleigh fading channels. Since Rayleigh fading channels are described in complex baseband, the SNR is defined as the effective per-dimension SNR, corresponding to the in-phase component of the received signal.
III-B Proof of Main Result
Based on the unified decoding rule (2), the decoding metric of CDF-ORBGRAND is given by
$\mathsf{D}(w)=\frac{1}{N}\sum_{i=1}^{N}\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\text{sgn}\left(\mathsf{T}_{i}\right)\cdot\mathsf{X}_{i}(w)<0\right),$ (6)
for $w=1,\cdots,\lceil e^{NR}\rceil$.
Since the terms in the summation in (6) involve rankings of channel reliability vector, they are correlated, and hence the standard GMI formula (see, e.g., [23, Eqn. (12)]) cannot be directly applied. Therefore we conduct the analysis directly from the first principle, akin to that in [10, 25]. We calculate the ensemble average error probability of decoding, which, for the i.i.d. random coding ensemble in our system model, is equivalent to the error probability of decoding when message $w=1$ is transmitted.
Under the condition that the transmitted message is $w=1$, some asymptotic properties of the CDF-ORBGRAND decoding metric (6) are derived and presented in the following lemmas. Their proofs are provided in the Appendix.
Lemma 1
As $N\rightarrow\infty$, the expectation of the decoding metric (6) for the transmitted codeword $w=1$ is given by
$\lim\limits_{N\to\infty}\mathbb{E}\mathsf{D}(1)$ $=\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y.$ (7)
Lemma 2
As $N\rightarrow\infty$, the variance of the decoding metric (6) for the transmitted codeword $w=1$ satisfies
$\lim\limits_{N\to\infty}\text{var}\mathsf{D}(1)=0.$ (8)
Lemma 3
As $N\rightarrow\infty$, for any non-transmitted codeword, i.e., $w^{\prime}\neq 1$, and for any $\theta<0$, the decoding metric (6) behaves almost surely as
$\lim\limits_{N\to\infty}\frac{1}{N}\ln\mathbb{E}\left\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\right\}$
$=$ $\mathbb{E}\Bigl[\ln\Bigl(1+e^{\theta\cdot\left\lvert\ln\frac{q^{+}(\mathsf{Y})}{q^{-}(\mathsf{Y})}\right\rvert}\Bigr)\Bigr]-\ln 2.$ (9)
With these lemmas, we proceed to prove Theorem 1.
For any $\epsilon>0$, define the event $\mathcal{U}_{\epsilon}$ as
$\mathcal{U}_{\epsilon}$ $=\Bigg\{\mathsf{D}(1)\geq\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\hskip 50.00008pt+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y+\epsilon\Bigg\},$
corresponding to the case where the decoding metric of the transmitted message, $\mathsf{D}(1)$, exceeds its expectation (Lemma 1) by at least $\epsilon$. Consequently, the ensemble average error probability of decoding can be expressed as
$\text{Pr}[\hat{\mathsf{W}}\neq 1]$
$=\quad$ $\text{Pr}[\hat{\mathsf{W}}\neq 1\lvert\mathcal{U}_{\epsilon}]\text{Pr}[\mathcal{U}_{\epsilon}]+\text{Pr}[\hat{\mathsf{W}}\neq 1,\mathcal{U}_{\epsilon}^{c}]$
$\leq\quad$ $\text{Pr}[\mathcal{U}_{\epsilon}]+\text{Pr}[\hat{\mathsf{W}}\neq 1,\mathcal{U}_{\epsilon}^{c}].$ (10)
Using Lemmas 1 and 2, along with Chebyshev inequality, we obtain that for any $\epsilon>0$,
$\lim\limits_{N\to\infty}$ $\text{Pr}\Bigg[\mathsf{D}(1)\geq\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y+\epsilon\Bigg]=0.$ (11)
Therefore, $\text{Pr}[\mathcal{U}_{\epsilon}]$ can be made arbitrarily small as $N$ tends to infinity.
On the other hand, applying the decoding rule (2) together with the union bound, we obtain
$\text{Pr}[\hat{\mathsf{W}}\neq 1,\mathcal{U}_{\epsilon}^{c}]$
$\leq\quad$ $\text{Pr}\Bigg[\exists w^{\prime}\neq 1,\mathsf{D}(w^{\prime})<\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y+\epsilon\Bigg]$
$\leq\quad$ $e^{NR}\text{Pr}\Bigg[\mathsf{D}(w^{\prime})<\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y+\epsilon\Bigg].$ (12)
Considering the probability in (12) conditioned with respect to $\underline{\mathsf{T}}$, and applying Chernoff bound, we obtain that for any $N$ and any $\theta<0$,
$-\frac{1}{N}\ln\text{Pr}\Bigg[\mathsf{D}(w^{\prime})<\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y+\epsilon\Big\lvert\underline{\mathsf{T}}\Bigg]$
$\geq\quad$ $\theta\Bigg[\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y+\epsilon\Bigg]$
$-\frac{1}{N}\ln\mathbb{E}\left\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\right\}.$ (13)
Letting $\epsilon\rightarrow 0,N\rightarrow\infty$, and applying the almost sure limit in Lemma 3, we obtain
$\text{Pr}\Bigg[\mathsf{D}(w^{\prime})<\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y+\epsilon\Big\lvert\underline{\mathsf{T}}\Bigg]$
[…]
$\hskip 40.00006pt+\theta\cdot\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\hskip 40.00006pt+\theta\cdot\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y\Bigg]\Bigg\}.$ (14)
Then applying the law of total expectation to remove the conditioning in the left side of (14) and substituting the right side upper bound into (12), we establish that the ensemble average error probability of decoding vanishes asymptotically as $N\rightarrow\infty$, whenever the rate $R$ satisfies
[…]
$\quad-\theta\cdot\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad-\theta\cdot\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y\Bigg\}.$ (15)
Now we prove that $I_{\text{CDF-ORB}}$ is equivalent to the mutual information $I(\mathsf{X};\mathsf{Y})$, i.e., symmetric capacity. By noting that $\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert=\ln\frac{q^{-}(y)}{q^{+}(y)}$ for $y\in\mathcal{R}_{1}$ and $\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert=\ln\frac{q^{+}(y)}{q^{-}(y)}$ for $y\in\mathcal{R}_{2}$, we can rewrite the mutual information (3) as
$I(\mathsf{X};\mathsf{Y})$ $=\ln 2-\frac{1}{2}\int_{\mathcal{R}_{1}}\ln\Bigl(1+e^{\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert}\Bigr)q^{+}(y)\mathrm{d}y$
$\quad-\frac{1}{2}\int_{\mathcal{R}_{2}}\ln\Bigl(1+e^{-\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert}\Bigr)q^{+}(y)\mathrm{d}y$
$\quad-\frac{1}{2}\int_{\mathcal{R}_{1}}\ln\Bigl(1+e^{-\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert}\Bigr)q^{-}(y)\mathrm{d}y$
$\quad-\frac{1}{2}\int_{\mathcal{R}_{2}}\ln\Bigl(1+e^{\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert}\Bigr)q^{-}(y)\mathrm{d}y.$ (16)
By using the fact that $q(y)=\frac{1}{2}(q^{+}(y)+q^{-}(y))$, we can further rewrite (16) as
[…]
$\quad-\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y-\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y.$ (17)
On the other hand, consider the right side of (15) as a function of $\theta<0$:
$I_{\text{CDF-ORB}}=\sup\limits_{\theta<0}f(\theta),$ (18)
[…]
$+\frac{\theta}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y+\frac{\theta}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y.$ (19)
Therefore, we obtain the following inequalities:
$I(\mathsf{X};\mathsf{Y})\geq I_{\text{CDF-ORB}}\geq f(-1)=I(\mathsf{X};\mathsf{Y}),$ (20)
implying that $I_{\text{CDF-ORB}}=I(\mathsf{X};\mathsf{Y})$, and that $f(\theta)$ attains its maximum at $\theta=-1$. This completes the proof.
IV Application of CDF-ORBGRAND in BICM
High-order modulation has been a prevailing technique in contemporary communication systems, for achieving high spectral efficiency. In order to apply GRAND to high-order modulation, we study BICM systems, where a high-order coded modulation channel is effectively decomposed into multiple bit channels, with the decoding metric for each bit channel separately computed. Remarkably, we establish that CDF-ORBGRAND is capable of achieving the BICM capacity, which was initially derived in the literature by treating the decomposed bit channels as independent.
IV-A BICM System Model
BICM employs an encoder, a bit interleaver $\pi$, and a binary labeling function $\mu:\left\{+1,-1\right\}^{m}\rightarrow\mathcal{S}$ that maps blocks of $m$ bits to symbols in a constellation $\mathcal{S}$. Let us start with a codebook of length $mN$ and rate $R$ nats per channel use, containing $\lceil e^{NR}\rceil$ codewords. When message $w$ is selected for transmission, the corresponding codeword is denoted by $\tilde{\underline{x}}(w)$. The codeword is then passed through the interleaver, and the resulting interleaved codeword is denoted by
$\underline{x}(w)=$ $\pi(\tilde{\underline{x}}(w))$
$=$ $[x_{1,1}(w),\cdots,x_{1,m}(w),\cdots,x_{N,1}(w),\cdots,x_{N,m}(w)].$
For notational convenience, $\underline{x}(w)$ is referred to as the codeword, and all subsequent descriptions are based on it. [^5] The original codeword $\tilde{\underline{x}}(w)$ can be recovered from $\underline{x}(w)$ through the de-interleaving operation. Since we consider random coding, elements of $\underline{x}(w)$ are i.i.d., uniformly drawn from $\{+1,-1\}$. After modulation, the $m$ bits $x_{i,1}(w),\ldots,x_{i,m}(w)$ as a block are mapped via the binary labeling function $\mu$ into $s_{i}(w)$, and the transmitted symbol sequence is denoted by
$\underline{s}(w)=[s_{1}(w),\cdots,s_{N}(w)].$
The elements of the codeword corresponding to the $j$-th bit of the labeling are denoted by
$\underline{x}_{j}(w)=[x_{1,j}(w),\cdots,x_{N,j}(w)],$
for $j=1,\ldots,m$. Conversely, we define the inverse mapping of the $j$-th bit of the labeling as $b_{j}:\mathcal{S}\rightarrow\left\{+1,-1\right\}$; that is, $b_{j}(s)$ is the codeword bit corresponding to the $j$-th position of the transmitted symbol $s$. The received sequence is denoted by $\underline{\mathsf{Y}}=[\mathsf{Y}_{1},\cdots,\mathsf{Y}_{N}]$.
We consider a general memoryless channel, where the output pdf of $\mathsf{Y}_{i}$ under input bit $x_{i,j}=+1$ is denoted by $q_{j}^{+}(y)$, and under input bit $x_{i,j}=-1$ is denoted by $q_{j}^{-}(y)$, respectively. These pdfs can be evaluated as
$q_{j}^{+}(y)=\frac{1}{\lvert\mathcal{S}_{j}^{+}\rvert}\sum\limits_{s\in\mathcal{S}_{j}^{+}}p(y\lvert s),\quad q_{j}^{-}(y)=\frac{1}{\lvert\mathcal{S}_{j}^{-}\rvert}\sum\limits_{s\in\mathcal{S}_{j}^{-}}p(y\lvert s),$
where $\mathcal{S}_{j}^{+}$ and $\mathcal{S}_{j}^{-}$ denote the sets of constellation symbols whose $j$-th bit equals +1+1 and 1-1, respectively. For each $j=1,\cdots,m$, We also assume that $q_{j}^{+}(y)$ and $q_{j}^{-}(y)$ satisfy Assumptions 1, 2 and 3 introduced in Section II-A.
Define the LLR $\mathsf{T}_{i,j}=\ln\frac{q_{j}^{+}(\mathsf{Y}_{i})}{q_{j}^{-}(\mathsf{Y}_{i})}$ and the reliability vector
$\lvert\mathsf{T}_{1,1}\rvert,\cdots,\lvert\mathsf{T}_{1,m}\rvert,\cdots,\lvert\mathsf{T}_{N,1}\rvert,\cdots,\lvert\mathsf{T}_{N,m}\rvert$].
For $i=1,\cdots,N$ and $j=1,\cdots,m$, denote by $\mathsf{R}_{i,j}$ the rank of $\lvert\mathsf{T}_{i,j}\rvert$ among the sorted array consisting of $\left\{\lvert\mathsf{T}_{1,1}\rvert,\cdots,\lvert\mathsf{T}_{1,m}\rvert,\cdots,\lvert\mathsf{T}_{N,1}\rvert,\cdots,\lvert\mathsf{T}_{N,m}\rvert\right\}$, from $1$ (the smallest) to $mN$ (the largest). Let $\Psi_{1}(t),\cdots,\Psi_{m}(t)$ be the CDFs of $\lvert\mathsf{T}_{i,1}\rvert,\cdots,\lvert\mathsf{T}_{i,m}\rvert$ for $t\geq 0$, respectively. We further define the average CDF as
$\bar{\Psi}(t)=\frac{1}{m}\sum_{j=1}^{m}\Psi_{j}(t).$ (21)
We also introduce the sets $\mathcal{R}_{1,j}=\left\{y\lvert q_{j}^{+}(y)<q_{j}^{-}(y)\right\}$ and $\mathcal{R}_{2,j}=\left\{y\lvert q_{j}^{+}(y)>q_{j}^{-}(y)\right\}$.
IV-B BICM Capacity
Analysis of BICM was initially conducted under the assumption of ideal interleaving [19], treating the BICM system model as $m$ mutually independent parallel binary-input memoryless channels, the $j$-th of which corresponds to the $j$-th bit of the labeling of the constellation symbols, with output pdf $q^{+}_{j}(y)$ under input $x=+1$ and $q^{-}_{j}(y)$ under input x=1x=-1. See Fig. 2 for an illustration.
Based on this parallel channel model, the BICM system achieves rate, i.e., the BICM capacity,
$C^{\text{BICM}}=\sum_{j=1}^{m}I(\mathsf{X}_{j};\mathsf{Y}),$ (22)
where in the $j$-th term of the summation, $\mathsf{X}_{j}$ is uniformly distributed over $\{+1,-1\}$, and $\mathsf{Y}$ is driven by $q^{\pm}_{j}(y)$, as described in the previous paragraph.
Encoder Binary-input channel 11 Binary-input channel 22 Binary-input channel $m$ $x\vdotsyy$
Figure 2: Parallel channel model of BICM with ideal interleaving.
Treating the $m$ bits within a symbol as independent results in a mismatched decoding rule, expressed as [26, 20]
$\hat{w}$ $=\mathop{\arg\max}\limits_{w=1,\cdots,\lceil e^{NR}\rceil}\prod_{i=1}^{N}\prod_{j=1}^{m}d_{j}(b_{j}(s_{i}(w)),y_{i}),$ (23)
where the $j$-th bit decoding metric is
$d_{j}(b_{j}(s)=+1,y)=q_{j}^{+}(y),d_{j}(b_{j}(s)=-1,y)=q_{j}^{-}(y).$
Via leveraging the tool of mismatched decoding, it has been shown that the BICM decoding rule (23) achieves the BICM capacity (22) even in the absence of ideal interleaving [21].
IV-C GRAND and Its Variants for BICM
After receiving a channel output vector $\underline{y}$, we compute the channel reliability vector
$\lvert t_{1,1}\rvert,\cdots,\lvert t_{1,m}\rvert,\cdots,\lvert t_{N,1}\rvert,\cdots,\lvert t_{N,m}\rvert$]
and the corresponding hard-decision vector
$\underline{x}_{\text{hard}}=[\kappa_{1,1},\cdots,\kappa_{1,m},\cdots,\kappa_{N,1},\cdots,\kappa_{N,m}],$
where $t_{i,j}=\ln\frac{q_{j}^{+}(y_{i})}{q_{j}^{-}(y_{i})}$ is the LLR and $\kappa_{i,j}=\text{sgn}(t_{i,j})$.
GRAND algorithms for BICM have a $2^{mN}\times mN$ $\pm 1$}-valued matrix $P$ for representing EPs, wherein the $q$-th row is denoted as
$P_{q}=[P_{q,1,1},\cdots,P_{q,1,m},\cdots,P_{q,N,1},\cdots,P_{q,N,m}],$
for conducting the $q$-th query: if $P_{q,i,j}=-1$, the sign of $\kappa_{i,j}$ in $\underline{x}_{\text{hard}}$ is flipped, and otherwise it remains unchanged. If $\underline{x}_{\text{hard}}$ after flipping is a codeword, it is declared as the decoder output; otherwise, if all rows of $P$ are exhausted without finding a codeword, a decoding failure is declared.
For ORBGRAND, we arrange the rows of $P$ so that the sum reliability of the $q$-th row, $\sum_{i,j:P_{q,i,j}=-1}r_{i,j}$, is non-decreasing with $q$, where $r_{i,j}$ is the rank of $\lvert t_{i,j}\rvert$. Similar to Section II-B, we have the following unified decoding rule for GRAND algorithms:
$\hat{w}=\mathop{\arg\min}\limits_{w=1,\cdots,\lceil e^{NR}\rceil}\frac{1}{mN}\sum_{i=1}^{N}\sum_{j=1}^{m}\gamma_{i,j}(\underline{y})\cdot\mathbf{1}(\kappa_{i,j}x_{i,j}(w)<0),$ (24)
for which different choices of $\gamma_{i,j}$ correspond to different variants of GRAND under BICM:
  • When $\gamma_{i,j}(\underline{y})=1$, (24) is the original GRAND;
  • When $\gamma_{i,j}(\underline{y})=\lvert t_{i,j}\rvert$, (24) is SGRAND, and is equivalent to the decoding rule in (23); see Appendix D for a proof of their equivalence;
  • When $\gamma_{i,j}(\underline{y})=\frac{r_{i,j}}{mN}$, (24) is ORBGRAND;
  • When $\gamma_{i,j}(\underline{y})=\bar{\Psi}^{-1}\left(\frac{r_{i,j}}{mN+1}\right)$, where $\bar{\Psi}^{-1}$ denotes the inverse function of $\bar{\Psi}$ given by (21), (24) is CDF-ORBGRAND.
IV-D Rate Analysis
The following theorem establishes that CDF-ORBGRAND achieves the BICM capacity.
Theorem 2
For the system model in Section IV-A, CDF-ORBGRAND achieves the BICM capacity, i.e.,
$I_{\text{CDF-ORB,BICM}}=C^{\text{BICM}}.$ (25)
Proof:
See Appendix E. ∎
For comparison, the following theorem provides the achievable rate of ORBGRAND for BICM.
Theorem 3
For the system model in Section IV-A, ORBGRAND achieves the GMI
$I_{\text{ORB,BICM}}$ $=m\ln 2-\inf\limits_{\theta<0}\Bigg\{\sum_{j=1}^{m}\Big(\int_{0}^{1}\ln(1+e^{\theta t})\mathrm{d}t$
$-\frac{\theta}{2}\int_{\mathcal{R}_{1,j}}\bar{\Psi}\Bigl(\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert\Bigr)q_{j}^{+}(y)\mathrm{d}y$
$-\frac{\theta}{2}\int_{\mathcal{R}_{2,j}}\bar{\Psi}\Bigl(\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert\Bigr)q_{j}^{-}(y)\mathrm{d}y\Big)\Bigg\}$ (26)
in nats/channel use.
Proof:
See Appendix E. ∎
IV-E Numerical Results
In this subsection, we numerically evaluate $I_{\text{CDF-ORB,BICM}}$ and $I_{\text{ORB,BICM}}$ for QPSK, 8PSK, and 16QAM, over Rayleigh fading channel with receiver-side perfect CSI. For each constellation, we consider both Gray and set-partitioning (SP) labeling schemes. As an illustration, Fig. 3 presents the 16QAM constellation diagrams corresponding to Gray and SP labelings. The channel input–output relationship is given by
$\mathsf{Y}=\mathsf{H}\mathsf{S}+\mathsf{Z},$
where $\mathsf{H}\sim\mathcal{CN}(0,1)$ is the channel fading coefficient, $\mathsf{S}$ is the channel input corresponding to a constellation symbol, and $\mathsf{Z}\sim\mathcal{CN}(0,1)$ is the AWGN. The average SNR is thus $\mathbb{E}[\lvert\mathsf{S}\rvert^{2}]$.
Figure 3: Gray (left) and SP (right) labelings for 16QAM. Here 0 and 11 correspond to 1-1 and +1+1, respectively.
The numerical results are shown in Fig. 4. It can be observed that ORBGRAND incurs a significant gap to the BICM capacity under both Gray and SP labelings. For example, with 8PSK and Gray labeling at SNR 15 dB, $I_{\text{ORB,BICM}}$ falls short of the BICM capacity (i.e., $I_{\text{CDF-ORB,BICM}}$) by approximately 0.5 bits per channel use. Moreover, this gap grows as the modulation order increases, reflecting the heightened sensitivity of ORBGRAND to high-order modulation in fading channels. Under the same conditions, the gap with SP labeling is even more pronounced than that with Gray labeling, indicating the impact of the choice of labeling scheme.
The performance gain of CDF-ORBGRAND over ORBGRAND can be attributed to the improved alignment of decoding metrics via rank companding, as commented at the beginning of Section III. This phenomenon appears to be more evident in the intermediate regime of SNR, and as modulation order increases.
In Fig. 4 we also include the achievable rates $\tilde{I}_{\text{ORB,BICM}}$ of BICM derived in [25] under the assumption of ideal interleaving; [^6] Similar to Fig. 1, here we rectify an error in the numerical program in [25] and update the numerical results. see the end of Appendix E for details. Although the derivation is approximate, we observe that the curves of $\tilde{I}_{\text{ORB,BICM}}$ are very close to the exact ones of ${I}_{\text{ORB,BICM}}$, indicating that the ideal interleaving assumption is usually valid in practice.
(a) Gray labeling
(b) Set-partitioning labeling
Figure 4: Achievable rates of ORBGRAND, CDF-ORBGRAND, and BICM capacity for QPSK, 8PSK, and 16QAM over Rayleigh fading channel with perfect CSI. Line styles indicate curve type: dashed for ORBGRAND with ideal interleaving, dash-dot for ORBGRAND with non-ideal interleaving, solid for CDF-ORBGRAND and BICM capacity. Line colors indicate modulation format: black for QPSK, blue for 8PSK, red for 16QAM.
V Conclusion
In this work, we first analyze the achievable rates of ORBGRAND and its variant under rank companding, CDF-ORBGRAND, from the perspective of the GMI for general binary-input memoryless channels. An somewhat unexpected yet fundamental finding is that the achievable rate of CDF-ORBGRAND is identical to the symmetric capacity, indicating that rank companding loses no information asymptotically. Building upon this finding, we extend the analysis to BICM and demonstrate that, CDF-ORBGRAND also achieves the BICM capacity. As numerical experiments reveal, ORBGRAND experiences a substantial rate gap in fading channels.
We emphasize that our analytical findings are information-theoretic in nature, which becomes effective as code lengths grow large. In practical scenarios with finite code lengths, CDF-ORBGRAND still exhibits a performance gap relative to ML decoding. Moreover, various practical optimization techniques for ORBGRAND have been proposed to enhance its performance in realistic settings; see, e.g., [7, 6, 10, 17, 18]. Exploring these methods and evaluating CDF-ORBGRAND under finite code lengths will be the focus of our future work.
Appendix A Proofs of Lemmas 1, 4, and 7
These three lemmas characterize the asymptotic behavior of the expectation of the decoding metric under $w=1$ as $N\rightarrow\infty$. We only provide the proof of Lemma 1. Lemmas 4 and 7 appear in the proofs of Theorems 2 and 3, respectively, in Appendix E. Their proofs are similar to that of Lemma 1 and are hence omitted.
For the decoding metric (6) of CDF-ORBGRAND, we have
$\mathbb{E}[\mathsf{D}(1)]=\frac{1}{N}\sum_{i=1}^{N}$
$\quad\mathbb{E}\Big[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\text{sgn}\left(\mathsf{T}_{i}\right)\cdot\mathsf{X}_{i}(1)<0\right)\Big].$ (27)
For each summand in (27), we have
$\mathbb{E}\Big[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\text{sgn}\left(\mathsf{T}_{i}\right)\cdot\mathsf{X}_{i}(1)<0\right)\Big]$
$=\quad$ $\frac{1}{2}\mathbb{E}\Big[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\mathsf{T}_{i}<0\right)\Big\lvert\mathsf{X}_{i}(1)=+1\Big]$
$+\frac{1}{2}\mathbb{E}\Big[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\mathsf{T}_{i}>0\right)\Big\lvert\mathsf{X}_{i}(1)=-1\Big].$ (28)
Considering the first term in (28) and applying the law of total expectation, we obtain
$\mathbb{E}\Big[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\mathsf{T}_{i}<0\right)\Big\lvert\mathsf{X}_{i}(1)=+1\Big]$
$=\quad$ $\mathbb{E}\Big[\mathbb{E}\Big[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\mathsf{T}_{i}<0\right)\Big\lvert\mathsf{X}_{i}(1)=+1,\mathsf{Y}_{i}\Big]\Big],$ (29)
for which
$\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\Big\lvert\mathsf{X}_{i}(1)=+1,\mathsf{Y}_{i}=y\Big]$
$=$ $\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\cdot\mathbf{1}(q^{+}(y)<q^{-}(y))\Big\lvert\mathsf{X}_{i}(1)=+1,\mathsf{Y}_{i}=y\Big]$
[…]
Note that $\mathsf{R}_{i}$ in (30) is the rank of $|t|=\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert$ when inserted into a sorted array of N1N-1 i.i.d. samples of $\lvert\mathsf{T}\rvert$, and we denote by $u(y)$ the upper branch of (30). Hence we have
$\mathbb{E}\Big[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\mathsf{T}_{i}<0\right)\Big\lvert\mathsf{X}_{i}(1)=+1\Big]$
$=\quad$ $\int_{\mathcal{R}_{1}}u(y)q^{+}(y)\mathrm{d}y.$ (31)
Similarly, the second term in (28) can be evaluated. So (28) is
$\mathbb{E}\Big[\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}\left(\text{sgn}\left(\mathsf{T}_{i}\right)\cdot\mathsf{X}_{i}(1)<0\right)\Big]$
$=\quad$ $\frac{1}{2}\int_{\mathcal{R}_{1}}u(y)q^{+}(y)\mathrm{d}y+\frac{1}{2}\int_{\mathcal{R}_{2}}u(y)q^{-}(y)\mathrm{d}y,$ (32)
which does not depend upon the index $i$, and consequently, (27) becomes
$\mathbb{E}[\mathsf{D}(1)]=\frac{1}{2}\int_{\mathcal{R}_{1}}u(y)q^{+}(y)\mathrm{d}y+\frac{1}{2}\int_{\mathcal{R}_{2}}u(y)q^{-}(y)\mathrm{d}y.$ (33)
The asymptotic behavior of $u(y)$ is given by Lemma 11 in Appendix F. We thus obtain
$\lim\limits_{N\to\infty}\mathbb{E}[\mathsf{D}(1)]$ $=\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y.$ (34)
Appendix B Proofs of Lemmas 2, 5, and 8
These three lemmas indicate that the variance of the decoding metric under $w=1$ asymptotically vanish as $N\rightarrow\infty$. Due to their similarity, we only provide the proof of Lemma 2. Lemmas 5 and 8 appear in the proofs of Theorems 2 and 3, respectively, in Appendix E.
Define $\mathsf{W}_{i}=\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\text{sgn}(\mathsf{T}_{i})\cdot\mathsf{X}_{i}(1)<0)$ and $\tilde{\mathsf{W}}_{i}=\mathsf{W}_{i}-\mathbb{E}\mathsf{W}_{i}$. Then for the decoding metric (6) of CDF-ORBGRAND, it follows that
$\text{var}\mathsf{D}(1)=\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathbb{E}[\tilde{\mathsf{W}}_{i}\tilde{\mathsf{W}}_{j}].$ (35)
For $i=j$, we have $\mathbb{E}[\tilde{\mathsf{W}}_{i}\tilde{\mathsf{W}}_{j}]=\mathbb{E}[\mathsf{W}_{i}^{2}]-\mathbb{E}[\mathsf{W}_{i}]^{2}$, where $\lim\limits_{N\to\infty}\mathbb{E}[\mathsf{W}_{i}]$ has been obtained in (34), and following similar steps as those in the proof of Lemma 1, we obtain
$\lim\limits_{N\to\infty}\mathbb{E}[\mathsf{W}_{i}^{2}]$ $=\frac{1}{2}\int_{\mathcal{R}_{1}}{\Big\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Big\rvert}^{2}q^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2}}{\Big\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Big\rvert}^{2}q^{-}(y)\mathrm{d}y.$ (36)
Based on Lemma 10 in Appendix F, we establish that $\lim\limits_{N\to\infty}\mathbb{E}[\tilde{\mathsf{W}}_{i}^{2}]$ remains finite. Since $\text{var}\mathsf{D}(1)$ contains $N$ such terms for $i=1,\ldots,N$, normalization by $N^{2}$ in (35) ensures that their contribution diminishes at a rate of $O\left(\frac{1}{N}\right)$ as $N$ grows without bound.
For $i\neq j$, we have $\mathbb{E}[\tilde{\mathsf{W}}_{i}\tilde{\mathsf{W}}_{j}]=\mathbb{E}[\mathsf{W}_{i}\mathsf{W}_{j}]-\mathbb{E}[\mathsf{W}_{i}]\mathbb{E}[\mathsf{W}_{j}]$. We now turn to the evaluation of $\mathbb{E}[\mathsf{W}_{i}\mathsf{W}_{j}]$, which, by definition, can be expressed as (37), shown at the top of this page. The first term in (37) leads to (38) and (39), shown at the top of this page and the next page, respectively. Note that $\mathsf{R}_{i}$ and $\mathsf{R}_{j}$ in (39) are the ranks resulting from inserting $|t_{i}|=\Bigl\lvert\ln\frac{q^{+}(y_{i})}{q^{-}(y_{i})}\Bigr\rvert$ and $|t_{j}|=\Bigl\lvert\ln\frac{q^{+}(y_{j})}{q^{-}(y_{j})}\Bigr\rvert$ into a sorted array of N2N-2 i.i.d. samples of $\lvert\mathsf{T}\rvert$, and we denote by $u(y_{i},y_{j})$ the upper branch of (39). Therefore, (38) can be rewritten as $\int_{\mathcal{R}_{1}}\int_{\mathcal{R}_{1}}u(y_{i},y_{j})q^{+}(y_{i})\mathrm{d}y_{i}q^{+}(y_{j})\mathrm{d}y_{j}$. Applying a similar procedure to the other three terms in (37), we obtain
$\mathbb{E}[\mathsf{W}_{i}\mathsf{W}_{j}]$ $=\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\text{sgn}(\mathsf{T}_{i})\cdot\mathsf{X}_{i}(1)<0)\mathbf{1}(\text{sgn}(\mathsf{T}_{j})\cdot\mathsf{X}_{j}(1)<0)\Big]$
$=\frac{1}{4}\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\mathbf{1}(\mathsf{T}_{j}<0)\Big\lvert\mathsf{X}_{i}(1)=+1,\mathsf{X}_{j}(1)=+1\Big]$
$\quad+\frac{1}{4}\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}>0)\mathbf{1}(\mathsf{T}_{j}<0)\Big\lvert\mathsf{X}_{i}(1)=-1,\mathsf{X}_{j}(1)=+1\Big]$
$\quad+\frac{1}{4}\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\mathbf{1}(\mathsf{T}_{j}>0)\Big\lvert\mathsf{X}_{i}(1)=+1,\mathsf{X}_{j}(1)=-1\Big]$
$\quad+\frac{1}{4}\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}>0)\mathbf{1}(\mathsf{T}_{j}>0)\Big\lvert\mathsf{X}_{i}(1)=-1,\mathsf{X}_{j}(1)=-1\Big],$ (37)
$\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\mathbf{1}(\mathsf{T}_{j}<0)\Big\lvert\mathsf{X}_{i}(1)=+1,\mathsf{X}_{j}(1)=+1\Big]$
$=\quad$ $\mathbb{E}\Big[\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\mathbf{1}(\mathsf{T}_{j}<0)\Big\lvert\mathsf{X}_{i}(1)=+1,\mathsf{X}_{j}(1)=+1,\mathsf{Y}_{i},\mathsf{Y}_{j}\Big]\Big],$ (38)
$\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\mathbf{1}(\mathsf{T}_{j}<0)\Big\lvert\mathsf{X}_{i}(1)=+1,\mathsf{X}_{j}(1)=+1,\mathsf{Y}_{i}=y_{i},\mathsf{Y}_{j}=y_{j}\Big]$
[…]
$\mathbb{E}[\mathsf{W}_{i}\mathsf{W}_{j}]$ $=\frac{1}{4}\int_{\mathcal{R}_{1}}\int_{\mathcal{R}_{1}}u(y_{i},y_{j})q^{+}(y_{i})\mathrm{d}y_{i}q^{+}(y_{j})\mathrm{d}y_{j}$
$\quad+\frac{1}{4}\int_{\mathcal{R}_{2}}\int_{\mathcal{R}_{1}}u(y_{i},y_{j})q^{+}(y_{i})\mathrm{d}y_{i}q^{-}(y_{j})\mathrm{d}y_{j}$
$\quad+\frac{1}{4}\int_{\mathcal{R}_{1}}\int_{\mathcal{R}_{2}}u(y_{i},y_{j})q^{-}(y_{i})\mathrm{d}y_{i}q^{+}(y_{j})\mathrm{d}y_{j}$
$\quad+\frac{1}{4}\int_{\mathcal{R}_{2}}\int_{\mathcal{R}_{2}}u(y_{i},y_{j})q^{-}(y_{i})\mathrm{d}y_{i}q^{-}(y_{j})\mathrm{d}y_{j}.$ (40)
By applying Lemma 12 in Appendix F, we obtain that $\lim\limits_{N\to\infty}\mathbb{E}[\tilde{\mathsf{W}}_{i}\tilde{\mathsf{W}}_{j}]=0$. Since $\text{var}\mathsf{D}(1)$ contains $N\times(N-1)$ such terms for $i,j=1,\ldots,N$, normalization by $N^{2}$ in (35) ensures that their contribution diminishes as $N$ grows without bound.
Summarizing the cases discussed, we see that $\text{var}\mathsf{D}(1)$ asymptotically vanishes as $N\rightarrow\infty$.
Appendix C Proofs of Lemmas 3, 6, and 9
C-A Proof of Lemma 3
For any $w^{\prime}\neq 1$, the conditional expectation of the CDF-ORBGRAND decoding metric (6) satisfies
$\mathbb{E}\left\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\right\}$
$=\quad$ $\mathbb{E}\left\{\prod_{i=1}^{N}e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\text{sgn}(\mathsf{T}_{i})\cdot\mathsf{X}_{i}(w^{\prime})<0)}\Big\lvert\underline{\mathsf{T}}\right\}$
$=\quad$ $\prod_{i=1}^{N}\mathbb{E}\left\{e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\text{sgn}(\mathsf{T}_{i})\cdot\mathsf{X}_{i}(w^{\prime})<0)}\Big\lvert\underline{\mathsf{T}}\right\},$ (41)
where we use the fact that $\underline{\mathsf{T}}$ is induced by $\underline{\mathsf{X}}(1)$ and is therefore independent of $\underline{\mathsf{X}}(w^{\prime})$. For each product term in (41), we have
$\mathbb{E}\left\{e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\text{sgn}(\mathsf{T}_{i})\cdot\mathsf{X}_{i}(w^{\prime})<0)}\Big\lvert\underline{\mathsf{T}}\right\}$
$=\quad$ $\frac{1}{2}\mathbb{E}\left\{e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\mathsf{T}_{i}<0)}\Big\lvert\underline{\mathsf{T}},\mathsf{X}_{i}(w^{\prime})=+1\right\}$
$+\frac{1}{2}\mathbb{E}\left\{e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\mathsf{T}_{i}>0)}\Big\lvert\underline{\mathsf{T}},\mathsf{X}_{i}(w^{\prime})=-1\right\}$
$=\quad$ $\frac{1}{2}e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\mathsf{T}_{i}<0)}+\frac{1}{2}e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\mathsf{T}_{i}>0)}$
$=\quad$ $\frac{1}{2}\Bigl(1+e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)}\Bigr),$ (42)
where since $\mathsf{R}_{i}$ is determined by $\underline{\mathsf{T}}$, the conditional expectation can be removed. Therefore, we have
$\frac{1}{N}\ln\mathbb{E}\Bigl\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\Bigr\}$
$=\quad$ $\frac{1}{N}\sum_{i=1}^{N}\ln\Bigl(1+e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)}\Bigr)-\ln 2$
$=\quad$ $\frac{1}{N}\sum_{n=1}^{N}\ln\Bigl(1+e^{\theta\Psi^{-1}\left(\frac{n}{N+1}\right)}\Bigr)-\ln 2,$ (43)
where we use the fact that $\left\{\mathsf{R}_{i}\right\}_{i=1\cdots,N}$ is a permutation of $\left\{1,\cdots,N\right\}$. Thus, we obtain
$\lim\limits_{N\to\infty}\frac{1}{N}\ln\mathbb{E}\Bigl\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\Bigr\}$
$=\quad$ $\lim\limits_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}\ln\Bigl(1+e^{\theta\Psi^{-1}\left(\frac{n}{N+1}\right)}\Bigr)-\ln 2$
$=\quad$ $\int_{0}^{1}\ln\Bigl(1+e^{\theta\Psi^{-1}(u)}\Bigr)\mathrm{d}u-\ln 2$
$=\quad$ $\int_{0}^{\infty}\ln\Bigl(1+e^{\theta t}\Bigr)\Psi^{\prime}(t)\mathrm{d}t-\ln 2$
$=\quad$ $\mathbb{E}\Bigl[\ln\Bigl(1+e^{\theta\cdot\Bigl\lvert\ln\frac{q^{+}(\mathsf{Y})}{q^{-}(\mathsf{Y})}\Bigr\rvert}\Bigr)\Bigr]-\ln 2,$ (44)
which completes the proof.
C-B Proof of Lemma 6
By following similar steps in the proof of Lemma 3, for any $w^{\prime}\neq 1$, the conditional expectation of the CDF-ORBGRAND decoding metric for BICM in (53) satisfies
$\frac{1}{N}\ln\mathbb{E}\Bigl\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\Bigr\}$
$=\quad$ $\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{m}\ln\Bigl(1+e^{\frac{\theta}{m}\bar{\Psi}^{-1}\left(\frac{\mathsf{R}_{i,j}}{mN+1}\right)}\Bigr)-m\ln 2$
$=\quad$ $\frac{1}{N}\sum_{n=1}^{mN}\ln\Bigl(1+e^{\frac{\theta}{m}\bar{\Psi}^{-1}\left(\frac{n}{mN+1}\right)}\Bigr)-m\ln 2.$ (45)
Thus, we obtain
$\lim\limits_{N\to\infty}\frac{1}{N}\ln\mathbb{E}\Bigl\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\Bigr\}$
$=\quad$ $m\int_{0}^{1}\ln\Bigl(1+e^{\frac{\theta}{m}\bar{\Psi}^{-1}(u)}\Bigr)\mathrm{d}u-m\ln 2$
$=\quad$ $m\int_{0}^{\infty}\ln\Bigl(1+e^{\frac{\theta}{m}t}\Bigr)\bar{\Psi}^{\prime}(t)\mathrm{d}t-m\ln 2$
$=\quad$ $m\int_{0}^{\infty}\ln\Bigl(1+e^{\frac{\theta}{m}t}\Bigr)\cdot\Bigl(\frac{1}{m}\sum_{j=1}^{m}\Psi_{j}^{\prime}(t)\Bigr)\mathrm{d}t-m\ln 2$
$=\quad$ $\sum_{j=1}^{m}\int_{0}^{\infty}\ln\Bigl(1+e^{\frac{\theta}{m}t}\Bigr)\Psi_{j}^{\prime}(t)\mathrm{d}t-m\ln 2$
$=\quad$ $\sum_{j=1}^{m}\mathbb{E}\Bigl[\ln\Bigl(1+e^{\frac{\theta}{m}\cdot\Bigl\lvert\ln\frac{q_{j}^{+}(\mathsf{Y})}{q_{j}^{-}(\mathsf{Y})}\Bigr\rvert}\Bigr)\Bigr]-m\ln 2,$ (46)
which completes the proof.
C-C Proof of Lemma 9
By following similar steps in the proof of Lemma 3, for any $w^{\prime}\neq 1$, the conditional expectation of the ORBGRAND decoding metric for BICM in (60) satisfies
$\frac{1}{N}\ln\mathbb{E}\Bigl\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\Bigr\}$ $=\frac{1}{N}\sum_{n=1}^{mN}\ln\Bigl(1+e^{\frac{\theta}{m}\cdot\frac{n}{mN}}\Bigr)-m\ln 2$
$\rightarrow m\int_{0}^{1}\ln(1+e^{\frac{\theta}{m}t})\mathrm{d}t-m\ln 2$ (47)
as $N\rightarrow\infty$, which completes the proof.
Appendix D Equivalence Between BICM Decoder (23) and SGRAND
The decoding rule of SGRAND for BICM can be equivalently expressed as
$\hat{w}$ $=\mathop{\arg\min}\limits_{w=1,\cdots,\lceil e^{NR}\rceil}\sum_{i=1}^{N}\sum_{j=1}^{m}$
$\quad\Bigl\lvert\ln\frac{q_{j}^{+}(y_{i})}{q_{j}^{-}(y_{i})}\Bigr\rvert\cdot\mathbf{1}\Bigl(\text{sgn}\Bigl(\ln\frac{q_{j}^{+}(y_{i})}{q_{j}^{-}(y_{i})}\Bigr)\cdot x_{i,j}(w)<0\Bigr).$ (48)
Likewise, the BICM decoding rule (23) can be equivalently expressed by
$\hat{w}$ $=\mathop{\arg\max}\limits_{w=1,\cdots,\lceil e^{NR}\rceil}\sum_{i=1}^{N}\sum_{j=1}^{m}\ln d_{j}(b_{j}(s_{i}(w)),y_{i}),$ (49)
where $d_{j}(+1,y)=q_{j}^{+}(y)$ and $d_{j}(-1,y)=q_{j}^{-}(y)$.
The summation in (48) can be expanded as
$\sum\limits_{(i,j):x_{i,j}(w)=+1,q_{j}^{+}(y_{i})<q_{j}^{-}(y_{i})}[\ln q_{j}^{-}(y_{i})-\ln q_{j}^{+}(y_{i})]$
$+$ $\sum\limits_{(i,j):x_{i,j}(w)=-1,q_{j}^{+}(y_{i})>q_{j}^{-}(y_{i})}[\ln q_{j}^{+}(y_{i})-\ln q_{j}^{-}(y_{i})],$ (50)
and the summation in (49) is
$\sum\limits_{(i,j):x_{i,j}(w)=+1}\ln q_{j}^{+}(y_{i})+\sum\limits_{(i,j):x_{i,j}(w)=-1}\ln q_{j}^{-}(y_{i}).$ (51)
The sum of (50) and (51), after simplification, turns out to be a constant
$\sum_{i=1}^{N}\sum_{j=1}^{m}\max\left\{\ln q_{j}^{+}(y_{i}),\ln q_{j}^{-}(y_{i})\right\}.$ (52)
Therefore, the two decoding rules are equivalent.
Appendix E Proofs of Theorems 2 and 3
E-A Proof of Theorem 2
Based on the unified decoding rule (24), the decoding metric of CDF-ORBGRAND for BICM is given by
$\mathsf{D}(w)$ $=\frac{1}{mN}\sum_{i=1}^{N}\sum_{j=1}^{m}\bar{\Psi}^{-1}\left(\frac{\mathsf{R}_{i,j}}{mN+1}\right)\cdot$
$\quad\mathbf{1}\left(\text{sgn}\left(\mathsf{T}_{i,j}\right)\cdot\mathsf{X}_{i,j}(w)<0\right),\quad w=1,\cdots,\lceil e^{NR}\rceil.$ (53)
Under the condition that the transmitted message is $w=1$, the following lemmas characterize some asymptotic properties of the CDF-ORBGRAND decoding metric (53). The proofs of these lemmas are provided in the preceding appendices.
Lemma 4
As $N\rightarrow\infty$, the expected value of the decoding metric (53) for the transmitted codeword is given by
$\lim\limits_{N\to\infty}\mathbb{E}\mathsf{D}(1)$ $=\frac{1}{m}\sum_{j=1}^{m}\Big(\frac{1}{2}\int_{\mathcal{R}_{1,j}}\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert q_{j}^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2,j}}\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert q_{j}^{-}(y)\mathrm{d}y\Big).$ (54)
Lemma 5
As $N\rightarrow\infty$, the variance of the decoding metric (53) for the transmitted codeword satisfies
$\lim\limits_{N\to\infty}\text{var}\mathsf{D}(1)=0.$ (55)
Lemma 6
As $N\rightarrow\infty$, for any non-transmitted codeword, i.e., $w^{\prime}\neq 1$, and for any $\theta<0$, the decoding metric (53) behaves almost surely as
$\lim\limits_{N\to\infty}\frac{1}{N}\ln\mathbb{E}\left\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\right\}$
$=\sum_{j=1}^{m}\mathbb{E}\Bigl[\ln\Bigl(1+e^{\frac{\theta}{m}\cdot\Bigl\lvert\ln\frac{q_{j}^{+}(\mathsf{Y})}{q_{j}^{-}(\mathsf{Y})}\Bigr\rvert}\Bigr)\Bigr]-m\ln 2.$ (56)
Similar to the proof of Theorem 1, we first define the event $\mathcal{U}_{\epsilon}$ that the decoding metric (53) for the transmitted codeword exceeds its asymptotic mean by at least $\epsilon$. This allows us to upper bound the ensemble average error probability of decoding by $\text{Pr}[\mathcal{U}_{\epsilon}]+\text{Pr}[\hat{w}\neq 1,\mathcal{U}_{\epsilon}^{c}]$. Using Lemmas 4 and 5 together with Chebyshev inequality, we obtain that $\text{Pr}[\mathcal{U}_{\epsilon}]\rightarrow 0$ as $N\rightarrow\infty$. On the other hand, by applying Lemma 6, we deduce that for any $R$ satisfying
$R$ $<I_{\text{CDF-ORB,BICM}}:=m\ln 2$
[…]
$\quad-\frac{\theta}{m}\cdot\frac{1}{2}\int_{\mathcal{R}_{1,j}}\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert q_{j}^{+}(y)\mathrm{d}y$
$\quad-\frac{\theta}{m}\cdot\frac{1}{2}\int_{\mathcal{R}_{2,j}}\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert q_{j}^{-}(y)\mathrm{d}y\Biggr)\Biggr\}$ (57)
in nats/channel use, the average error probability of decoding asymptotically vanishes as $N\to\infty$.
By setting $\eta=\frac{\theta}{m}$, the right side of (57) can be written as
[…]
$+\frac{\eta}{2}\int_{\mathcal{R}_{1,j}}\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert q_{j}^{+}(y)\mathrm{d}y+\frac{\eta}{2}\int_{\mathcal{R}_{2,j}}\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert q_{j}^{-}(y)\mathrm{d}y\Biggr)\Biggr\}$
$=\sup\limits_{\eta<0}\Bigg\{\sum_{j=1}^{m}f_{j}(\eta)\Bigg\}.$ (58)
From the proof of Theorem 1 in Section III-B, we know that for each $j=1,\ldots,m$, $f_{j}(\eta)$ attains its extreme at $\eta=-1$ and $f_{j}(-1)=I(\mathsf{X}_{j};\mathsf{Y})$. Hence, we have the achievable rate of CDF-ORBGRAND for BICM
$I_{\text{CDF-ORB,BICM}}=\sup\limits_{\eta<0}\Big\{\sum_{j=1}^{m}f_{j}(\eta)\Big\}=\sum_{j=1}^{m}I(\mathsf{X}_{j};\mathsf{Y})=C^{\text{BICM}},$ (59)
thereby completing the proof.
E-B Proof of Theorem 3
Based on the unified decoding rule (24), the decoding metric of ORBGRAND for BICM is given by
$\mathsf{D}(w)$ $=\frac{1}{mN}\sum_{i=1}^{N}\sum_{j=1}^{m}\frac{\mathsf{R}_{i,j}}{mN}\cdot$
$\quad\mathbf{1}\left(\text{sgn}\left(\mathsf{T}_{i,j}\right)\cdot\mathsf{X}_{i,j}(w)<0\right),w=1,\cdots,\lceil e^{NR}\rceil.$ (60)
Under the condition that the transmitted message is $w=1$, the following lemmas characterize some asymptotic properties of the ORBGRAND decoding metric (60). The proofs of these lemmas are provided in the preceding appendices.
Lemma 7
As $N\rightarrow\infty$, the expected value of the decoding metric (60) for the transmitted codeword is given by
$\lim\limits_{N\to\infty}\mathbb{E}\mathsf{D}(1)$ $=\frac{1}{m}\sum_{j=1}^{m}\Big(\frac{1}{2}\int_{\mathcal{R}_{1,j}}\bar{\Psi}\Bigl(\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert\Bigr)q_{j}^{+}(y)\mathrm{d}y$
$\quad+\frac{1}{2}\int_{\mathcal{R}_{2,j}}\bar{\Psi}\Bigl(\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert\Bigr)q_{j}^{-}(y)\mathrm{d}y\Big).$ (61)
Lemma 8
As $N\rightarrow\infty$, the variance of the decoding metric (60) for the transmitted codeword satisfies
$\lim\limits_{N\to\infty}\text{var}\mathsf{D}(1)=0.$ (62)
Lemma 9
As $N\rightarrow\infty$, for any non-transmitted codeword, i.e., $w^{\prime}\neq 1$, and for any $\theta<0$, the decoding metric (60) behaves almost surely as
$\lim\limits_{N\to\infty}\frac{1}{N}\ln\mathbb{E}\left\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\right\}$
$=m\int_{0}^{1}\ln(1+e^{\frac{\theta}{m}t})\mathrm{d}t-m\ln 2.$ (63)
Based on these lemmas, we can complete the proof of Theorem 3 following steps similar to those used in the proofs of Theorems 1 and 2.
As a side remark, we compare ${I}_{\text{ORB,BICM}}$ with the achievable rate derived in our previous work [25], where we assumed ideal interleaving (see Fig. 2) and hence the achievable rate is simply the sum of GMIs over the $m$ bit channels, i.e.,
$\tilde{I}_{\text{ORB,BICM}}$ $=\sum_{j=1}^{m}I_{\text{ORB},j}$
$=m\ln 2-\sum_{j=1}^{m}\inf\limits_{\theta_{j}<0}\Bigg\{\int_{0}^{1}\ln(1+e^{\theta_{j}t})\mathrm{d}t$
$\quad-\theta_{j}\cdot\frac{1}{2}\int_{\mathcal{R}_{1,j}}\Psi_{j}\Bigl(\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert\Bigr)q_{j}^{+}(y)\mathrm{d}y$
$\quad-\theta_{j}\cdot\frac{1}{2}\int_{\mathcal{R}_{2,j}}\Psi_{j}\Bigl(\Bigl\lvert\ln\frac{q_{j}^{+}(y)}{q_{j}^{-}(y)}\Bigr\rvert\Bigr)q_{j}^{-}(y)\mathrm{d}y\Bigg\}.$ (64)
In contrast to ${I}_{\text{ORB,BICM}}$ given by (26), here the expression (64) suggests that coding is done separately over different bit channels, rather than across them jointly. Only when all bit channels have identical laws, i.e., $q_{j}^{\pm}(y)$ being the same for $j=1,\ldots,m$, ${I}_{\text{ORB,BICM}}$ and $\tilde{I}_{\text{ORB,BICM}}$ coincide. Nevertheless, as shown in our numerical results in Section IV-E, their discrepancy is typically negligible.
Appendix F Supporting Lemmas
Lemma 10
If $q^{+}(y)$ and $q^{-}(y)$ satisfy Assumptions 1 and 2, then
$\frac{1}{2}\int_{\mathcal{R}_{1}}{\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert}^{2}q^{+}(y)\mathrm{d}y+\frac{1}{2}\int_{\mathcal{R}_{2}}{\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert}^{2}q^{-}(y)\mathrm{d}y$
$-\Bigl(\frac{1}{2}\int_{\mathcal{R}_{1}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{+}(y)\mathrm{d}y+\frac{1}{2}\int_{\mathcal{R}_{2}}\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert q^{-}(y)\mathrm{d}y\Bigr)^{2}$ (65)
is a finite value.
Proof:
Considering the first term in (65), there exists a sufficiently large constant $M$ such that
$\int_{\mathcal{R}_{1}}{\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert}^{2}q^{+}(y)\mathrm{d}y$ (66)
$\leq$ $\int_{\lvert y\rvert<M}{\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert}^{2}q^{+}(y)\mathrm{d}y+\int_{\lvert y\rvert\geq M}{\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert}^{2}q^{+}(y)\mathrm{d}y$
$\leq$ $\int_{\lvert y\rvert<M}{\Bigl\lvert\ln\frac{q^{+}(y)}{q^{-}(y)}\Bigr\rvert}^{2}q^{+}(y)\mathrm{d}y+\int_{\lvert y\rvert\geq M}S_{2}^{2}(\lvert y\rvert)S_{1}(\lvert y\rvert)e^{-a\lvert y\rvert}\mathrm{d}y.$ (67)
The first term of (67) is the integral of a bounded function over a bounded interval, so it is finite; the second term of (67) is also finite, as light-tailed distributions admit all moments. Therefore, (66) is bounded. Applying analogous argument to the other three terms in (65) shows that all the integrals involved are finite, and hence (65) is finite. ∎
Lemma 11
Suppose that the CDFs of $\lvert\mathsf{T}_{\cdot,1}\rvert,\ldots,\lvert\mathsf{T}_{\cdot,m}\rvert$ all satisfy Assumption 3. Denote by $\mathsf{R}(v)$ the rank of $v$ when inserted into the sorted array consisting of independent samples $\left\{\lvert\mathsf{T}_{1,1}\rvert,\cdots,\lvert\mathsf{T}_{1,m}\rvert,\cdots,\lvert\mathsf{T}_{N,1}\rvert,\cdots,\lvert\mathsf{T}_{N,m}\rvert\right\}$, with an arbitrary one of them removed. We have
$\lim\limits_{N\to\infty}\frac{\mathbb{E}[\mathsf{R}(v)]}{mN}=\bar{\Psi}(v),$ (68)
$\lim\limits_{N\to\infty}\frac{\mathbb{E}[{\mathsf{R}(v)}^{2}]}{m^{2}N^{2}}=\bar{\Psi}(v)^{2},$ (69)
$\lim\limits_{N\to\infty}\mathbb{E}\Big[\bar{\Psi}^{-1}\Big(\frac{\mathsf{R}(v)}{mN+1}\Big)\Big]=v,$ (70)
$\lim\limits_{N\to\infty}\mathbb{E}\Big[\Bigl(\bar{\Psi}^{-1}\Big(\frac{\mathsf{R}(v)}{mN+1}\Big)\Bigr)^{2}\Big]=v^{2}.$ (71)
Proof:
Without loss of generality, we remove $\lvert\mathsf{T}_{N,m}\rvert$, and the proof holds with very minor modification if we remove any other sample.
Based on the definition of $\mathsf{R}(v)$, we have
$\mathsf{R}(v)=1+\sum_{i=1}^{N-1}\sum_{j=1}^{m}\mathbf{1}(\lvert\mathsf{T}_{i,j}\rvert<v)+\sum_{j=1}^{m-1}\mathbf{1}(\lvert\mathsf{T}_{N,j}\rvert<v).$ (72)
Taking the expectation of (72) yields
$\mathbb{E}[\mathsf{R}(v)]=1+(N-1)\sum_{j=1}^{m}\Psi_{j}(v)+\sum_{j=1}^{m-1}\Psi_{j}(v),$ (73)
from which (68) immediately follows. Taking the expectation of the square of (72) and conducting a few algebraic manipulation steps yield (69).
Denoting $G(\cdot)=\bar{\Psi}^{-1}(\cdot)$, $\mathsf{S}=\frac{\mathsf{R}(v)}{mN+1}$ and $p=\bar{\Psi}(v)$, we can show that as $N\rightarrow\infty$,
$\mathbb{E}[\mathsf{S}-p]=O\Bigl(\frac{1}{N}\Bigr),\mathbb{E}[(\mathsf{S}-p)^{2}]=O\Bigl(\frac{1}{N}\Bigr),$ (74)
$\mathbb{E}[\lvert\mathsf{S}-p\rvert^{3}]\leq O\Bigl(\frac{1}{N^{\frac{3}{2}}}\Bigr).$ (75)
We perform a Taylor expansion of $G(\mathsf{S})$ around $p$:
$G(\mathsf{S})=G(p)+G^{\prime}(p)(\mathsf{S}-p)+\frac{1}{2}G^{\prime\prime}(p)(\mathsf{S}-p)^{2}+K_{3},$ (76)
where the residual term satisfies
$\lvert K_{3}\rvert\leq\frac{1}{6}\sup\limits_{u\in[p-\delta,p+\delta]}\lvert G^{\prime\prime\prime}(u)\rvert\cdot\lvert\mathsf{S}-p\rvert^{3}.$
Taking the expectation of (76), substituting (74) and (75), and applying Assumption 3, we obtain (70).
Similarly, performing a Taylor expansion of $G(\mathsf{S})^{2}$ around $p$ and taking its expectation, we obtain (71). ∎
Lemma 12
Suppose that the CDFs of $\lvert\mathsf{T}_{\cdot,1}\rvert,\ldots,\lvert\mathsf{T}_{\cdot,m}\rvert$ all satisfy Assumption 3. Denote by $\mathsf{R}(v_{a})$ and $\mathsf{R}(v_{b})$ the ranks of $v_{a}$ and $v_{b}$, respectively, when inserted into the sorted array consisting of independent samples $\left\{\lvert\mathsf{T}_{1,1}\rvert,\cdots,\lvert\mathsf{T}_{1,m}\rvert,\cdots,\lvert\mathsf{T}_{N,1}\rvert,\cdots,\lvert\mathsf{T}_{N,m}\rvert\right\}$, with arbitrary two of them removed. We have
$\lim\limits_{N\to\infty}\frac{\mathbb{E}[\mathsf{R}(v_{a})\mathsf{R}(v_{b})]}{m^{2}N^{2}}=\bar{\Psi}(v_{a})\bar{\Psi}(v_{b}),$ (77)
$\lim\limits_{N\to\infty}\mathbb{E}\Big[\bar{\Psi}^{-1}\Big(\frac{\mathsf{R}(v_{a})}{mN+1}\Big)\bar{\Psi}^{-1}\Big(\frac{\mathsf{R}(v_{b})}{mN+1}\Big)\Big]=v_{a}v_{b}.$ (78)
Proof:
Without loss of generality, we remove $\lvert\mathsf{T}_{N,m-1}\rvert$ and $\lvert\mathsf{T}_{N,m}\rvert$, and the proof holds with very minor modification if we remove any other two samples.
Based on the definition of $\mathsf{R}(v_{a})$ and $\mathsf{R}(v_{b})$, we can deduce that
$\mathbb{E}[\mathsf{R}(v_{a})\mathsf{R}(v_{b})]=N^{2}\sum_{i=1}^{m}\sum_{j=1}^{m}\Psi_{i}(v_{a})\Psi_{j}(v_{b})+O(N),$ (79)
from which (77) immediately follows.
Denoting $G(\cdot)=\bar{\Psi}^{-1}(\cdot),\mathsf{S}_{a}=\frac{\mathsf{R}(v_{a})}{mN+1},\mathsf{S}_{b}=\frac{\mathsf{R}(v_{b})}{mN+1},p_{a}=\bar{\Psi}(v_{a})$ and $p_{b}=\bar{\Psi}(v_{b})$, and following the similar approach for the proof of (70), we perform Taylor expansions of $G(\mathsf{S}_{a})$ and $G(\mathsf{S}_{b})$ around $p_{a}$ and $p_{b}$, respectively; then multiplying these Taylor expansions and taking the expectation lead to (78). ∎